Please show Every Step and As Soon as Possible.Protective layer A 25-mm diameter, spherical fuel element (27.6 W/m·K) with a volumetric rate of generation of 15 x 106 W/m²
Answer #1

Question |Ts;=314-15°C Solution : Tsi Tso q* = 154106 w m3 Qg=Q Stead eady Condition Oq= eneqy generation Rate. Rate across PInside Raduis of Protective Layer is R2 Ri=9=Radius of Fuel element Riz 12.5 mm = = 12.5x10m Outside Radues of Poorective LayQg 9 122.718 = Tsi-Tso 1.2575205 Psi-Tso = 154.32 Tso = Ts;: -154-32 Tso = 314.15 - 154.32 Tso 159.83 °c Ans